If g(x) is the inverse function of f(x) and f`(x) =1/1+x^4 then g`(x) is

Grade:12

1 Answers

Kishori dube
5 years ago
Given: g(x) is invese of function f(x)Then f(g(x))=xDifferentiating both side wrt x,we getf`(g(x))*g`(x)=1So g`(x)=1/f`(g(x)) =1+[g(x)]^4